OpenStudy (flor):

scheduling the games for softball league to find the no. of games he needs to schedule , he uses the equation 1/2n2-1/2n where a represents the number of games needed for each team exactly once and n represents the no. of teams. write the equation in factored form. how many games are needed for 7 teams to plan each other exactly 3 times

OpenStudy (anonymous):

\[= \frac{1}{2n^2} - \frac{1}{2n} = \frac{(1- n)}{2n^2}\]

OpenStudy (anonymous):

it must be \[\frac{1}{2}n^2-\frac{1}{2}n\] right?

OpenStudy (anonymous):

or even \[\frac{n(n-1)}{2}\]

OpenStudy (anonymous):

\[\frac{1}{2}n^2-\frac{1}{2}n = 9\]

OpenStudy (anonymous):

\[n^2 - n = 18\]

OpenStudy (anonymous):

something goofy with this question. what does "where 9 represents the number of games needed for each team exactly once " mean?

OpenStudy (flor):

yes that is